OpenStudy (anonymous):

Richard had $52 before he went shopping. Richard bought 3 shirts of equal price. He had $9 left. Which equation could be used to determine how much each shirt cost (x)? A. 52 + 9x = 3 B. 52 + 9 = 3x C. 52 - 3x = 9 D. 52 - 3 = 9x

OpenStudy (anonymous):

had 52- the three shirts at x left with 9

OpenStudy (apoorvk):

let price of each shirt be $x. so 3 shirts cost 3x dollars. so money left is initial - final so, 52-3x=9

OpenStudy (anonymous):

C. 52 - 3x = 9 since 3x + 9 = 52